TGBT19A_502_01 - Additional LBW Astronomical Commissioning


  • Test all incoherent LBW modes in cal, search, and fold
  • Get fluxcal data in all LBW modes
  • Do on-sky tests of all spectral line FS modes to compare with data from TGBT17B_501


  • Session begins at 04:45. Joe is switching to the LBW branch.
  • Switch over took a while, but scheduling block submitted at 05:24
  • i0100x0512
    • Bandpass looks good
    • No cal signal seen in cal scan.
    • Still seeing some additional smearing and off-pulse noise
  • i0100x1024
    • Bandpass levels are a bit low but acceptable. Could safely increase by a factor of 2.0
    • No issues in cal scan
    • Pulsar looks good at first glance in fold mode
    • Search mode seems to have a frequency shift again. VEGAS input IF is 350 MHz, as it should be. Still not clear what is causing this. There was no frequency shift in the total-intensity scan.
  • i0100x2048
    • Levels too high in scan #9. Redoing.
    • Scan #10 looks better with vegas.scale = 1000. Even that could be lowered a bit.
    • There is a frequency shift in this mode as well. IF tuning is all correct. Will try to recreate it if time at the end.
  • i0100x4096
    • No obvious frequency shift.
  • i0200x1024
    • Cal scan ran without issue.
    • Abort at start of fold-mode scan, but it seemed to come from the antenna manager. This may have been caused by the operator starting and stopping some weather-related programs in Task Master
    • No issue with subsequent scans
  • i0200x2048
    • Forgot to change vegas.conf to use the new BOF file with the fixed 10-GbE core. I also had to modify vegas.conf so that the packet format for this mode was 1SFA, which it appears to be based on packet size.
    • After making the above changes scans ran fine, but there was a frequency shift in this mode, similar to the i0100x2048 mode.
  • i0200x4096
    • Bandpass looks good
    • Pulsar detected. No apparent frequency shift.
  • i0200x8192
    • Bandpass looks good
    • Pulsar detected. No apparent frequency shift.
  • Going to try to recreate the frequency shift issues and then run those modes in a loop to see if they persist.
    • Sequence starting with scan #23
    • No apparent shift in any of the i0100x1024 scans.
    • Also no visible shift in any of the i0100x2048 scans.
    • No visible shift in i0200x2048, either
    • Could not recreate any frequency shifts
  • Moving on to fluxcal scans on 3C43
    • c0100x0064 scans didn't run properly. All others were fine.
    • Redid c0100x0064 fluxcal scans at end.
  • Moving on to spectral line checkout scans
    • First scan in sequence is #65
    • Oops. Forgot to have the operator retract the boom. Then had some VNC issues. Lost a bit of time.
    • Good scans start with #74
  • Also got position-switched data in Modes 1, 3, 4, 7, 10, 19, 20, and 29. Also 2, 5, 6, 8, 9, 11, 12, 13, 14, 15, 16, 17, 18, 21, 22, 23
  • Joe began switching back to release branch at 12:45.
  • Performed checkout scans. All looks good.


  • LBW modes all seem to check out except for the intermittent frequency shifts in i0100x1024, i0100x2048, and i0200x2048. These can't be reliably reproduced and will still require more investigation.
  • Ryan will need to modify auto-plotting script to work in LBW modes when not all banks are writing data.
  • Overall a successful run. Barring any unforeseen issues with the final processing I think we will be ready to release.
-- RyanLynch - 2019-04-13
Topic revision: r6 - 2019-04-14, RyanLynch
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback